# On the Convergence Rates of Federated Q-Learning across Heterogeneous Environments

### v1

@@ Abstract
Large-scale multi-agent systems are often deployed across wide geographic areas, where agents interact with heterogeneous environments. There is an emerging interest in understanding the role of heterogeneity in the performance of the federated versions of classic reinforcement learning algorithms. In this paper, we study synchronous federated Q-learning, which aims to learn an optimal Q-function by having $K$ agents average their local Q-estimates per $E$ iterations. We observe an interesting phenomenon on the convergence speeds in terms of $K$ and $E$. Similar to the homogeneous environment settings, there is a linear speed-up concerning $K$ in reducing the errors that arise from sampling randomness. Yet, in sharp contrast to the homogeneous settings, $E>1$ leads to significant performance degradation. The slow convergence of having $E>1$ turns out to be fundamental rather than an artifact of our analysis. We prove that, for a wide range of stepsizes, the $\ell_{\infty}$ norm of the error cannot decay faster than $\Theta(E/T)$. In addition, our experiments demonstrate that the convergence exhibits an interesting two-phase phenomenon. For any given stepsize, there is a sharp phase-transition of the convergence: the error decays rapidly in the beginning yet later bounces up and stabilizes. Provided that the phase-transition time can be estimated, choosing different stepsizes for the two phases leads to faster overall convergence.
@@ 1 Introduction
Advancements in unmanned capabilities are rapidly transforming industries and national security by enabling fast-paced and versatile operations across domains such as advanced manufacturing (Park et al., 2019), autonomous driving (Kiran et al., 2021), and battlefields (Möhlenhof et al., 2021). Reinforcement learning (RL) – a cornerstone for unmanned capabilities – is a powerful machine learning method that aims to enable an agent to learn an optimal policy via interacting with its operating environment to solve sequential decision-making problems (Bertsekas & Tsitsiklis, 1996; Bertsekas, 2019). However, the ever-increasing complexity of the environment results in a high-dimensional state-action space, imposing overwhelmingly high sample collection requirements on individual agents. This limited-data challenge becomes a significant hurdle that must be addressed to realize the potential of reinforcement learning.
In this paper, we study reinforcement learning within a federated learning framework (also known as Federated Reinforcement Learning (Qi et al., 2021; Jin et al., 2022; Woo et al., 2023)), wherein multiple agents independently collect samples and collaboratively train a common policy under the orchestration of a parameter server without disclosing the local data trajectories. A simple illustration can be found in Fig. 1. When the environments of all agents are homogeneous, it has been shown that the federated version of classic reinforcement learning algorithms can significantly alleviate the data collection burden on individual agents (Woo et al., 2023; Khodadadian et al., 2022) – error bounds derived therein exhibit a linear speedup in terms of number of agents. Moreover, by tuning the synchronization period $E$ (i.e., the number of iterations between agent synchronization), the communication cost can be significantly reduced compared with $E=1$ yet without significant performance degradation.
Figure 1: A federated learning system instance.
However, many large-scale multi-agent systems are often deployed across wide geographic areas, resulting in agents interacting with heterogeneous environments. For instance, connected and autonomous vehicles (CAVs) operating in various regions of a metropolitan area encounter diverse conditions such as varying traffic patterns, road infrastructure, and local regulations. The clients’ federation must be managed to ensure the robustness of the learned policy against environmental heterogeneity.
@@
Contributions. In this paper, we study synchronous federated Q-learning with environment heterogeneity.
  • We provide a fine-grained characterization of the error evolution, which decays to zero as the number of iterations $T$ increases. We observe an interesting phenomenon on the convergence speeds in terms of $K$ and $E$, where $K$ is the number of agents. Similar to the homogeneous environment settings, there is a linear speed-up concerning $K$ in reducing the errors that arise from sampling randomness. Yet, in sharp contrast to the homogeneous settings, $E>1$ leads to significant performance degradation.
  • We prove that the convergence slowing down for $E>1$ is fundamental. We show that the $\ell_{\infty}$ norm of the error cannot decay faster than $\Theta(E/T)$. A practical implication of this impossibility result is that, eventually, having multiple local updates (i.e., $E>1$) ends up consuming more samples (i.e., $E\times$ more) than using $E=1$.
  • Our numerical results illustrate that when the environments are heterogeneous and $E>1$, and there exists a sharp phase-transition of the error convergence: The error decays rapidly in the beginning yet later bounces up and stabilizes. In addition, provided that the phase-transition time can be estimated, choosing different stepsizes for the two phases can lead to faster overall convergence.
@@ 2 Related Work
FedAvg was proposed in the seminal work (McMahan et al., 2017), and has been one of the most widely implemented federated learning algorithms. It also has inspired many follow-up algorithms such as FedProx (Li et al., 2020b), FedNova (Wang et al., 2020), SCAFFOLD (Karimireddy et al., 2020), and adaptive federated methods (Deng et al., 2020). Despite intensive efforts, the theoretical understanding of FedAvg is far from complete. Most existing theoretical work on FedAvg overlooks the underlying data statistics at the agents, which often leads to misalignment of the pessimistic theoretical predictions and empirical success (Su et al., 2023; Pathak & Wainwright, 2020; Wang et al., 2022a, b). This theory and practice gap is studied in a recent work (Su et al., 2023) when solving general non-parametric regression problems. It shows that the limiting points of the global model under FedAvg is one unbiased estimator of the underlying model that generates the data.
@@
Reinforcement Learning. There has been extensive research on the convergence guarantees of reinforcement learning algorithms. A recent surge of work focuses on non-asymptotic convergence and the corresponding sample complexity for the single-agent setup. Bhandari et al. (2018) analyses non-asymptotic TD learning with linear function approximation (LFA) considering a variety of noise conditions, including noiseless, independent noise and Markovian noise. The results were extended to TD($\lambda$) and Q-learning. Li et al. (2020a) investigates the sample complexity of asynchronous Q-learning with different families of learning rates. They also provide an extension of using variance reduction methods inspired by the seminal SVRG algorithm. Li et al. (2024) shows the sample complexity of Q-learning. Let ${\mathcal{A}}$ be the set of actions. When $|{\mathcal{A}}|=1$, the sample complexity of synchronous Q-learning is sharp and minimax optimal, however, when $|{\mathcal{A}}|\geq 2$, it is shown that synchronous Q- learning has a lower bound which is not minimax optimal.
@@
Federated Reinforcement Learning. Woo et al. (2023) provides sample complexity guarantees for both synchronous and asynchronous distributed Q-learning and reveals that given the same transition probability (i.e., homogeneous environment) for all agents, they can speed up the convergence process linearly by collaboratively learning the optimal Q-function. Doan et al. (2019) investigates distributed Temporal Difference (TD) algorithm TD(0) with LFA under the setting of multi-agent MDP, where multiple agents act in a shared environment and each agent has its own reward function. They provide a finite-time analysis of this algorithm that with constant stepsize, the estimates of agents can converge to a neighborhood around optimal solutions at the rate of ${\mathcal{O}}(1/T)$ and asymptotically converge to the optimal solutions at the rate of ${\mathcal{O}}(1/\sqrt{T+1})$, where $T$ is the timestep. Khodadadian et al. (2022) studies on-policy federated TD learning, off-policy federated TD learning, and federated Q-learning of homogeneous environment and reward with Markovian noise. The sample complexity derived exhibits linear speedup with respect to the number of agents. Heterogeneous environments are considered in Jin et al. (2022); Wang et al. (2023); Xie & Song (2023); Zhang et al. (2023b). Jin et al. (2022) studies federated Q-learning and policy gradient methods under the setting of different known transition probabilities for each agent. Yet, no state sampling is considered. Wang et al. (2023) proposes FedTD(0) with LFA dealing with the environmental and reward heterogeneity of MDPs. They rigorously prove that in a low-heterogeneity regime, there is a linear convergence speedup in the number of agents. Xie & Song (2023) uses KL-divergence to penalize the deviation of local update from the global policy, and they prove that under the setting of heterogeneous environments, the local update is beneficial for global convergence using their method. Zhang et al. (2023a) proposes FedSARSA using the classic on-policy RL algorithm SARSA with linear function approximation (LFA) under the setting of heterogeneous environments and rewards. They theoretically prove that the algorithm can converge to the near-optimal solution. Neither Xie & Song (2023) nor Zhang et al. (2023a) characterize sample complexity.
@@ 3 Preliminary on Q-Learning
$V^{\pi}(s)=\mathbb{E}_{a_{t}\sim\pi(\cdot|s_{t}),s_{t+1}\sim P(\cdot|s_{t},a_{t})}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\mid s_{0}=s\right],\quad\forall s\in{\mathcal{S}}.$
The state-action value function, or Q-function $Q^{\pi}:{\mathcal{S}}\times{\mathcal{A}}\rightarrow\mathbb{R}$, evaluates the expected total discounted reward from taking action $a$ in state $s$ and then following policy $\pi$:
$Q^{\pi}(s,a)=R(s,a)\,+\,\mathbb{E}_{a_{t}\sim\pi(\cdot|s_{t}),s_{t+1}\sim P(\cdot|s_{t},a_{t})}\left[\sum_{t=1}^{\infty}\gamma^{t}{\mathcal{R}}(s_{t},a_{t})\mid s_{0}=s,a_{0}=a\right],\,\forall(s,a)\in{\mathcal{S}}\times{\mathcal{A}}.$
@@
where $s^{\prime}$ is sampled from the environment or the transition probability and $\lambda$ is the stepsize.
@@ 4 Federated Q-learning
The federated learning system consists of one parameter server (PS) and $K$ agents. The $K$ agents are deployed in possibly heterogeneous yet independent environments. The $K$ agents are modeled as Markov Decision Processes (MDPs) with ${\mathcal{M}}_{k}=\langle{\mathcal{S}},{\mathcal{A}},{\mathcal{P}}^{k},\gamma,R\rangle$ for $k=1,\cdots,K$, where ${\mathcal{P}}^{k}=\{P^{k}(\cdot\mid s,a)\}_{s\in{\mathcal{S}},a\in{\mathcal{A}}}$ are the collection of probability distributions that can be heterogeneous across agents. In the synchronous setting, each agent $k$ has access to a generative model, and generates a new state sample for each $(s,a)$ pair via
$s_{t}^{k}(s,a)\sim P^{k}(\cdot\mid s,a)\phantom{==}$
i.e., $\mathbb{P}\left\{s_{t}^{k}(s,a)=s^{\prime}\right\}=P^{k}(s^{\prime}\mid s,a)$ for all $s^{\prime}\in{\mathcal{S}}$, independently across state-action pairs $(s,a)$. For each $(s,a)$, the transition dynamics of the global environment $\bar{P}(\cdot\mid s,a)$ (Jin et al., 2022) is defined as
$\vspace{-1em}\bar{P}(s^{\prime}\mid s,a)=\frac{1}{K}\sum_{k=1}^{K}P^{k}(s^{\prime}\mid s,a),\forall s^{\prime}.$ (1)
@@
Let $Q^{*}$ denote the optimal Q-function of the global MDP. By the Bellman optimality equation, we have for all $(s,a)$,
$Q^{*}(s,a)=R(s,a)+\gamma\sum_{s^{\prime}\in{\mathcal{S}}}\bar{P}(s^{\prime}\mid s,a)V^{*}(s^{\prime}),$ (3)
@@
7:         $Q_{t+\frac{1}{2}}^{k}(s,a)=(1-\lambda)Q_{t}^{k}(s,a)+\lambda\left(R(s,a)+\gamma\max_{a^{\prime}\in{\mathcal{A}}}\,Q_{t}^{k}(s_{t}^{k}(s,a),a^{\prime})\right)$.
@@
The goal of the federated Q-learning is to have the $K$ agents collaboratively learn $Q^{*}$. We consider synchronous federated Q-learning, which is a natural integration of FedAvg and Q-learning (Woo et al., 2023; Jin et al., 2022) – described in Algorithm 1. Every agent initializes its local $Q^{k}$ estimate as $Q_{0}$ and performs standard synchronous Q-learning based on the locally collected samples $s_{t}^{k}(s,a)$. Whenever $t+1\mod E=0$, through the parameter server, the $K$ agents average their local estimate of $Q$; that is, all agents report their $Q_{t+\frac{1}{2}}^{k}$ to the parameter server, which computes the average and sends back to agents.

### v2

@@ Abstract
Large-scale multi-agent systems are often deployed across wide geographic areas, where agents interact with heterogeneous environments. There is an emerging interest in understanding the role of heterogeneity in the performance of the federated versions of classic reinforcement learning algorithms. In this paper, we study synchronous federated Q-learning, which aims to learn an optimal Q-function by having $K$ agents average their local Q-estimates per $E$ iterations. We provide a characterization of the error evolution, which decays to zero as the number of iterations $T$ increases. We show that when K(E1)K(E-1) is below a certain threshold, similar to the homogeneous environment settings, there is a linear speed-up concerning $K$. In sharp contrast, when K(E1)K(E-1) is above the threshold, heterogeneous environments lead to significant performance degradation. In particular, as $E$ increases, the convergence rate deteriorates. The slow convergence of having $E>1$ turns out to be fundamental rather than an artifact of our analysis. We prove that, for a wide range of stepsizes, the $\ell_{\infty}$ norm of the error cannot decay faster than $\Theta_{R}\left(E/{((1-\gamma)T)}\right)$, where $\Theta_{R}$ only hides numerical constants and the specific choice of reward values. In addition, our experiments demonstrate that the convergence exhibits an interesting two-phase phenomenon. For any given stepsize, there is a sharp phase transition of the convergence: the error decays rapidly in the beginning yet later bounces up and stabilizes.
@@ 1 Introduction
Advancements in unmanned capabilities are rapidly transforming industries and national security by enabling fast-paced and versatile operations across domains such as advanced manufacturing (Park et al., 2019), autonomous driving (Kiran et al., 2021), and battlefields (Möhlenhof et al., 2021). Reinforcement learning (RL) – a cornerstone for unmanned capabilities – is a powerful machine learning method that aims to enable an agent to learn an optimal policy via interacting with its operating environment to solve sequential decision-making problems (Bertsekas & Tsitsiklis, 1996; Bertsekas, 2019). However, the ever-increasing complexity of the environment results in a high-dimensional state-action space, often imposing overwhelmingly high sample collection requirements on individual agents. This limited-data challenge becomes a significant hurdle that must be addressed to realize the potential of reinforcement learning.
In this paper, we study reinforcement learning within a federated learning framework (also known as Federated Reinforcement Learning (Qi et al., 2021; Jin et al., 2022; Woo et al., 2023)), wherein multiple agents independently collect samples and collaboratively train a common policy under the orchestration of a parameter server without disclosing the local data trajectories. A simple illustration can be found in Fig. 1. When the environments of all agents are homogeneous, it has been shown that the federated version of classic reinforcement learning algorithms can significantly alleviate the data collection burden on individual agents (Woo et al., 2023; Khodadadian et al., 2022) – error bounds derived therein exhibit a linear speedup in terms of the number of agents. Moreover, by tuning the synchronization period $E$ (i.e., the number of iterations between agent synchronization), the communication cost can be significantly reduced compared with $E=1$
Figure 1: An illustration of a federated learning system.
yet without significant performance degradation. However, many large-scale multi-agent systems are often deployed across wide geographic areas, resulting in agents interacting with heterogeneous environments. For instance, connected and autonomous vehicles (CAVs) operating in various regions of a metropolitan area encounter diverse conditions such as varying traffic patterns, road infrastructure, and local regulations. Intuitively, the environmental heterogeneity may lead to misaligned learning signals across agents, potentially hinder the convergence, and degrade the generalization performance of the learned policies. Hence, the clients’ federation must be managed in a way that ensures the learned policy is robust to environmental heterogeneity.
@@
Contributions. Our contributions can be summarized as follows. All the asymptotic notations, e.g., $\mathcal{O}$} and $\widetilde{\mathcal{O}}$, unless otherwise specified, hide only numerical constants.
  • We characterize the error evolution of synchronous Federated Q-learning, showing that it decays to zero as the number of iterations $T$ increases. When K(E1)K(E-1) is below a threshold of $\widetilde{{\mathcal{O}}}\left({(\kappa\epsilon)^{-1}(1-\gamma)^{-2}}\right)$, similar to the homogeneous environment settings, there is a linear speed-up concerning $K$ and the sample complexity is $\widetilde{\mathcal{O}}\left(\frac{|{\mathcal{S}}||{\mathcal{A}}|}{K(1-\gamma)^{5}\epsilon^{2}}\right)$, matching the homogenous setting (Woo et al., 2023). Here, $\mathcal{S}$ and $\mathcal{A}$ are the state and action spaces, $\gamma\in(0,1)$ denotes the discount factor, and $\kappa$ is a scalar characterizing the environment heterogeneity. In sharp contrast, when K(E1)K(E-1) is above the threshold, heterogeneous environments lead to significant performance degradation and results in a unique sample complexity of $\widetilde{\mathcal{O}}\left(\frac{|\mathcal{S}||\mathcal{A}|\kappa E}{(1-\gamma)^{3}\epsilon}\right)$. Note that $E$ is also $\epsilon$-dependent. Hence, the sample complexity $\widetilde{\mathcal{O}}\left(\frac{|\mathcal{S}||\mathcal{A}|\kappa E}{(1-\gamma)^{3}\epsilon}\right)$ does not contradict the lower bound of $\Omega(\epsilon^{-2})$ for the single-agent case.
  • We prove that the convergence slowing down for $E>1$ is fundamental. We show that the $\ell_{\infty}$ norm of the error cannot decay faster than $\Theta_{R}\left(\frac{E}{(1-\gamma)T}\right)$, where $\Theta_{R}$ only hides numerical constants and the specific choice of reward values. A practical implication of this impossibility result is that, eventually, having multiple local updates (i.e., $E>1$) ends up consuming more samples (i.e., $E\times$ more) than using $E=1$ to reach a target accuracy.
  • Our numerical results illustrate that when the environments are heterogeneous and $E>1$, there exists a sharp phase-transition of the error convergence for not too small stepsizes: The error decays rapidly in the beginning yet later bounces up and stabilizes. In addition, provided that the phase-transition time can be estimated, choosing different stepsizes for the two phases can lead to faster overall convergence for both constant and time-decaying stepsizes. We conjecture that this is because the error in phase 1 is mainly controlled by the initial error with impacting factor decay exponentially in time, and the error in phase 2 is dominated by the collective perturbation caused by environment heterogeneity and multiple local updates (i.e., $E>1$).
@@ 2 Related Work
FedAvg was proposed in the seminal work (McMahan et al., 2017), and has been one of the most widely implemented federated learning algorithms. It also has inspired many follow-up algorithms such as FedProx (Li et al., 2020b), FedNova (Wang et al., 2020), SCAFFOLD (Karimireddy et al., 2020), and adaptive federated methods (Deng et al., 2020). Despite intensive efforts, the theoretical understanding of FedAvg is far from complete. Most existing theoretical work on FedAvg overlooks the underlying data statistics at the agents, which often leads to misalignment of the pessimistic theoretical predictions and empirical success (Su et al., 2023; Pathak & Wainwright, 2020; Wang et al., 2022a; b). This theory and practice gap was studied in a recent work (Su et al., 2023) in the context of solving general non-parametric regression problems.
@@
Reinforcement Learning. For the single-agent setup, there has been extensive research on the convergence guarantees of reinforcement learning algorithms. A recent surge of work studied non-asymptotic convergence and the corresponding sample complexity. Bhandari et al. (2018) analyzed non-asymptotic Temporal Difference (TD) learning with linear function approximation under a variety of noise conditions, including noiseless, independent noise, and Markovian noise. The results were extended to TD($\lambda$) and Q-learning. Li et al. (2020a) investigated the sample complexity of asynchronous Q-learning with different families of learning rates. They also provided an extension of using variance reduction methods inspired by the seminal SVRG algorithm. Li et al. (2024) shows the sample complexity of Q-learning. Recall that $\mathcal{A}$} is the set of actions. When $|{\mathcal{A}}|=1$, the sample complexity of synchronous Q-learning is sharp and minimax optimal; however, when $|{\mathcal{A}}|\geq 2$, it was shown that synchronous Q-learning has a lower bound which is not minimax optimal.
Multi-Agent RL. Yu et al. (2022) tested multi-agent Proximal Policy Optimization in four multi-agent testbeds wherein agents fully share the parameters, and showed its competitive performance. Christianos et al. (2021) proposed a selective parameter sharing technique, which automatically partitions agents so that they can benefit from the parameter sharing. Zhong et al. (2024) further proposed provably correct heterogeneous-agent algorithms, which allow agents to have different policy functions. The algorithms showed superior effectiveness and stability in various challenging benchmarks.
Federated Reinforcement Learning. Woo et al. (2023) provided sample complexity guarantees for both synchronous and asynchronous distributed Q-learning. They revealed that, under the same transition probability (i.e., homogeneous environment) for all agents, the convergence speed in learning the optimal Q-function can be accelerated linearly in the number of agents. They also uncovered the blessing of heterogeneity in terms of state-action exploration – a completely different notion of heterogeneity from our focus. Salgia & Chi (2025) explored the frontier of sample and communication complexities under homogeneous environments. Via variance reduction and communication quantization, they designed an algorithm that achieves order-optimal sample and communication complexities. Doan et al. (2019) investigated the distributed TD(0) with linear function approximation for a setting where multiple agents act in a shared environment and each agent has its own reward function. Khodadadian et al. (2022) studied on-policy federated TD learning, off-policy federated TD learning, and federated Q-learning of homogeneous environments and reward with Markovian noises. The sample complexity derived exhibits linear speedup with respect to the number of agents.
@@
Heterogeneous environments were considered in Jin et al. (2022); Wang et al. (2023); Xie & Song (2023); Zhang et al. (2023). Jin et al. (2022) studied federated Q-learning and policy gradient methods assuming known transition probabilities. To address heterogeneity in both environments and rewards, Wang et al. (2023) proposed FedTD(0) with linear function approximation. They proved that, in a low-heterogeneity regime, there is a linear convergence speedup in the number of agents. Xie & Song (2023) used KL-divergence to penalize the deviation of local update from the global policy, and proved that the local update is beneficial for global convergence. Zhang et al. (2024) proposed FedSARSA using the classic on-policy RL algorithm SARSA with linear function approximation. They theoretically proved that the algorithm can converge to a near-optimal solution. Neither Xie & Song (2023) nor Zhang et al. (2024) characterized sample complexity.
Technical comparisons with Woo et al. (2023); Zhang et al. (2024); Wang et al. (2023).
Zhang et al. (2024) and Wang et al. (2023) examined federated versions of TD learning and SARSA, whereas our paper studied federated Q-learning, which offers distinct theoretical and practical advantages for optimal policy learning. Specifically, the upper bound in Zhang et al. (2024) and Wang et al. (2023) do not indicate how fundamentally the convergence rates are impacted by the heterogeneity $\kappa$ and synchronization period $E$. In addition, their upper bounds do not decay to 0 as $T\rightarrow\infty$. Our upper bound converges to 0 as $T\rightarrow\infty$. Furthermore, we derived a lower bound on the convergence rates, showing the fundamental limitation of multiple local updates (i.e., $E>1$) in the presence of environmental heterogeneity. To the best of our knowledge, this is the first result of its kind.
While our analysis of Theorem 1 builds upon the roadmap established by Woo et al. (2023), adapting their analysis to our setting introduces significant challenges. In Woo et al. (2023), agents operate in homogeneous environments, i.e., each of the $K$ agents shares the same transition distributions. This homogeneity allows the concentration bound on the difference between the true transition distribution and sampled estimates to become arbitrarily small as the number of samples increases. However, in our setting, each agent has its own environment with a distinct transition distribution. This heterogeneity introduces a perturbation term in the error upper bound that does not decrease with additional samples. Additionally, when $E>1$ and $\kappa>0$, the term involving $\kappa(E-1)$ in the upper bound becomes the dominant term, resulting in a unique sample complexity. Further technical details and implications of these adjustments are provided in Corollary 2.
Work RL
Algorithm
Hetero-
geneity
Optimality Lower bound Sampling Finite-time Task
Wang et al. (2023) TD(0) Pred
Xie & Song (2023) Policy Gradient Pred, Plan
Zhang et al. (2024) SARSA Pred, Plan
Khodadadian et al. (2022) TD,
Q-Learning
Pred, Plan
Jin et al. (2022) Q-Learning,
Policy Gradient
Pred, Plan
Woo et al. (2023) Q-Learning Pred, Plan
Zheng et al. (2023) Q-Learning Pred, Plan
Our work Q-Learning Pred, Plan
Table 1: Comparison of various works in the context of FRL. Pred and Plan stand for prediction (policy evaluation) and planning (policy optimization), respectively.
@@ 3 Preliminary on Q-Learning
$V^{\pi}(s)=\mathbb{E}_{a_{t}\sim\pi(\cdot|s_{t}),s_{t+1}\sim P(\cdot|s_{t},a_{t})}\left[\sum_{t}\gamma^{t}R(s_{t},a_{t})\mid s_{0}=s\right],\quad\forall s\in{\mathcal{S}}.$
The Q-function (a.k.a. state-action value function), $Q^{\pi}:{\mathcal{S}}\times{\mathcal{A}}\rightarrow\mathbb{R}$, evaluates the expected total discounted reward from taking action $a$ in state $s$ and then following policy $\pi$:
$Q^{\pi}(s,a)=R(s,a)\,+\,\mathbb{E}_{a_{t}\sim\pi(\cdot|s_{t}),s_{t+1}\sim P(\cdot|s_{t},a_{t})}\left[\sum_{t}\gamma^{t}R(s_{t},a_{t})\mid s_{0}=s,a_{0}=a\right],\,\forall(s,a)\in{\mathcal{S}}\times{\mathcal{A}}.$
@@
where $s^{\prime}\sim P(\cdot\mid s,a)$, and $\lambda$ is the stepsize.
@@ 4 Federated Q-learning
The federated learning system consists of one parameter server (PS) and $K$ agents. The $K$ agents are deployed in possibly heterogeneous yet independent environments. The $K$ agents are modeled as MDPs with ${\mathcal{M}}_{k}=\langle{\mathcal{S}},{\mathcal{A}},{\mathcal{P}}^{k},\gamma,R\rangle$ for $k=1,\cdots,K$, where ${\mathcal{P}}^{k}=\{P^{k}(\cdot\mid s,a)\}_{s\in{\mathcal{S}},a\in{\mathcal{A}}}$ is the collection of probability distributions. In the synchronous setting, each agent $k$ has access to a generative model. At each iteration $t$, it generates a new state sample $s_{t}^{k}(s,a)\sim P^{k}(\cdot\mid s,a)$ for each $(s,a)$, i.e., $\mathbb{P}\left\{s_{t}^{k}(s,a)=s^{\prime}\right\}=P^{k}(s^{\prime}\mid s,a)$ for all $s^{\prime}\in{\mathcal{S}}$, independently across state-action pairs $(s,a)$. For each $(s,a)$, the global environment $\bar{P}(\cdot\mid s,a)$ (Jin et al., 2022) is defined as
$\vskip-10.00002pt\bar{P}(s^{\prime}\mid s,a)=\frac{1}{K}\sum_{k=1}^{K}P^{k}(s^{\prime}\mid s,a),\forall s^{\prime}$ (1)
@@
Let $Q^{*}$ denote the optimal Q-function of the global MDP. By the Bellman optimality equation, we have,
$Q^{*}(s,a)=R(s,a)+\gamma\sum_{s^{\prime}\in{\mathcal{S}}}\bar{P}(s^{\prime}\mid s,a)V^{*}(s^{\prime}),~~\forall(s,a)$ (3)
@@
7:   $Q_{t+\frac{1}{2}}^{k}(s,a)=(1-\lambda)Q_{t}^{k}(s,a)+\lambda\left(R(s,a)+\gamma\max_{a^{\prime}\in{\mathcal{A}}}\,Q_{t}^{k}(s_{t}^{k}(s,a),a^{\prime})\right)$
@@
The goal of federated Q-learning is to have the $K$ agents collaboratively learn $Q^{*}$. We consider synchronous federated Q-learning, which is a natural integration of FedAvg and Q-learning (Woo et al., 2023; Jin et al., 2022) – described in Algorithm 1. Every agent initializes its local $Q^{k}$ estimate as $Q_{0}$ and performs standard synchronous Q-learning based on the locally collected samples $s_{t}^{k}(s,a)$. Whenever $t+1\mod E=0$, through the parameter server, the $K$ agents average their local estimate of $Q$; that is, all agents report their $Q_{t+\frac{1}{2}}^{k}$ to the parameter server, which computes the average and sends back to agents.